# Dissect-and-Restore: AI-based Code Verification with Transient Refactoring

### v1

@@ 2.1. Formal specification challenges
$2.$ Verifier-specific challenges
@@
There are even more unexpected cases that may require a disproportionate amount of time to fix. For instance, proving that $|A|<|B|$ when AA is a subset of $B$ appears to be trivial, however, a proof in Dafny requires to manually assert |AB|==0|A-B|==0. These quirks derive from the fact that verifiers translate code to low level SMT formulas and feed these formulas into a solver, which is unaware of the original abstractions and semantics used by the verifiers.
@@ 3. Prometheus: Transiently Simplifying Code
The proof decomposition (§3.3) tackles the challenges outlined in Sec. 2.3. Prometheus breaks down complex proof obligations into a hierarchy/tree of structured helper lemmas and sub-lemmas. These lemmas capture inductive steps or auxiliary properties that are generally difficult for LLMs to infer. The goal of this decomposition is to enable LLMs to focus on simpler, localized proof sub-goals, aligning with the insights of Sec. $2.$ Proof decomposition also plays a key role in addressing challenges introduced by formal specification misalignments (Sec. 2.1). By isolating verification into smaller proof goals, Prometheus makes it possible to bridge these gaps incrementally, without requiring the specification to be rewritten.
@@ 3.2. Code-level decomposition
Sub-method verification. Each sub-method is verified sequentially by the proof-level decomposition described in Sec. $3.$ When Dafny verification fails, the cause often falls into one of two categories: $(i)$ an insufficient postcondition in a called method e.g., the postcondition of innerPost may be too weak to support the verification of MaxSubImpl (e.g., ensures true), or $(ii)$ missing intermediate proof steps within the current method. In the former case, Prometheus prompts the LLM to strengthen the callee’s contract; in the latter, the system proceeds with further proof-level decomposition within the current method.
@@
$3.$ Proof-level decomposition
@@ 4. Evaluation
LLMs selection. We primarily use Claude Sonnet 3.7 (Anthropic, 2025), as the Claude series models have demonstrated superior performance in generating Dafny code (Loughridge et al., 2024). Claude Sonnet 3.7 was the latest available version when this work began, ensuring it was not exposed to our experiments or prompts. Additionally, we utilize OpenAI o4-mini (OpenAI, 2025) for checking code verifiability, as it excels in complex reasoning at the time of submission. All the models run with the following configuration: initial temperature of $0.5$, the maximum number of tokens is $4028$, and the timeout is 2020 s. We selected these values to strike a balance between creativity, concise responses, and reasonable generation time.
Evaluation configuration. To prevent infinite verification loops, we set Dafny’s verification timeout to 2020 seconds. Within Prometheus, each node of the tree is allowed a maximum of $t=10$ generation attempts before rolling back. The regeneration counter $s$ is set to 22, and $\Delta{temp}$ to $0.3$. This means that each node has two generation attempts with temperatures of $0.5$ and $0.2$. Finally, we set a hard timeout of $500$ seconds on the generation process for both the baseline and Prometheus, after which the attempt is aborted.
@@ 4.1. Verification Performance (Q1)
Prometheus achieves higher success rate on tasks involving nested loops and complex reasoning. We begin by demonstrating Prometheus’s ability to tackle non-trivial verification tasks using the TitanBench benchmark. As shown in Table 1, our system successfully completes 1919 out of $22$ tasks, achieving a success rate of 86%, compared to the baseline’s 68%. Among the $22$ tasks, $12$ involve programs with at least one nested loop. Importantly, the four additional successful verifications by Prometheus come from these nested loop tasks. While the baseline solves 66 out of $10$ such tasks, Prometheus successfully solves all of them. Upon reviewing the solutions provided by Prometheus for these four tasks, we found that verifying them requires either $(i)$ multiple invariants for each loop, or $(ii)$ additional helper lemmas beyond a single method.
@@ 4.3. Robustness across formal specifications (Q3)
Prometheus effectively handles specification-code misalignment, especially in the presence of nested loops. Table 3 shows the results of the experiment, with tasks categorized by the presence/absence of nested loops. Among the 13 tasks, the baseline successfully verifies the 4 simplest ones. Notably, these tasks do not involve any arithmetic operations over the input array, which significantly simplifies verification. Without any arithmetic aggregation across the elements of the array, the inner loop does not need to maintain or reason about an evolving state, making the proof much more tractable. However, the baseline struggles significantly with tasks that include nested loops and aggregate state, aligning with our observations in Sec. 2.1 and Sec. 2.4 that AI-based systems tend to fail when multiple reasoning steps are required and to deal with misaligned, yet more readable, formal specifications. Conversely, Prometheus substantially outperforms the baseline by verifying 99 among the $13$ tasks, including those with nested loops. Prometheus’ code decomposition mechanism enables it to modularize loops effectively, while its proof decomposition component introduces the necessary helper lemmas. Together, these features allow Prometheus to verify 44 tasks that involve nested loops. In the three failed cases, Prometheus succeeds in proposing helper lemmas to bridge the misalignment, but fails in proving them. These lemmas are not trivial and require more advanced reasoning. We leave addressing this challenge for future work.
@@
$4.$ Scaling to Complex Programs (Q4)
@@
The selected examples offer a balanced and meaningful range of complexity, with task difficulty determined by the number of verification lines and the number of helper lemmas required in the manually verified code. As shown in Table 4, the length of the verification code (LOVC) ranges from 2121 to $487$ lines, and the number of helper lemmas ranges from 11 to $11$. Since the algorithms are more challenging to verify, we increase the generation timeout to $1500$ seconds. Because the complexity of these verification tasks surpasses the reasoning capacity of today’s LLMs, we supply a proof outline to both the baseline and Prometheus. [^4] The textual proofs are available in the artifact. These outlines ensure that the verification process begins from a sound and purposeful foundation, and they highlight how human insights can effectively complement LLM-based verification.
Prometheus can prove algorithms that require >>200 lines of verification code. Looking at the results in Table 4, the baseline is able to verify only the two simplest tasks, while Prometheus successfully verifies 77 tasks within five runs. Interestingly, Prometheus often provides more lemmas than the corresponding human-written solutions, meaning that the system tends to decompose proofs into smaller, more manageable steps, while humans can deal with complex reasoning within a single, more comprehensive lemma. In the unsolved Task #8, Prometheus is capable of solving partial proof obligations but fails to generate all of them. We note that Prometheus starts to fail more often as the proof context grows in complexity. Nevertheless, the partially verified program can still be manually inspected by the user, who may complete the verification starting from the well-defined, partially correct proof.

### v2

@@ 2.1. Formal specification challenges
2.2. Verifier-specific challenges
@@
There are even more unexpected cases that may require a disproportionate amount of time to fix. For instance, proving that $|A|<|B|$ when $A$ is a subset of $B$ appears to be trivial, however, a proof in Dafny requires to manually assert |AB|==0|A-B|==0. These quirks derive from the fact that verifiers translate code to low level SMT formulas and feed these formulas into a solver, which is unaware of the original abstractions and semantics used by the verifiers.
@@ 3. Prometheus: Transiently Simplifying Code
The proof decomposition (§3.3) tackles the challenges outlined in Sec. 2.3. Prometheus breaks down complex proof obligations into a hierarchy/tree of structured helper lemmas and sub-lemmas. These lemmas capture inductive steps or auxiliary properties that are generally difficult for LLMs to infer. The goal of this decomposition is to enable LLMs to focus on simpler, localized proof sub-goals, aligning with the insights of Sec. 2.2. Proof decomposition also plays a key role in addressing challenges introduced by formal specification misalignments (Sec. 2.1). By isolating verification into smaller proof goals, Prometheus makes it possible to bridge these gaps incrementally, without requiring the specification to be rewritten.
@@ 3.2. Code-level decomposition
Sub-method verification. Each sub-method is verified sequentially by the proof-level decomposition described in Sec. 3.3. When Dafny verification fails, the cause often falls into one of two categories: $(i)$ an insufficient postcondition in a called method e.g., the postcondition of innerPost may be too weak to support the verification of MaxSubImpl (e.g., ensures true), or $(ii)$ missing intermediate proof steps within the current method. In the former case, Prometheus prompts the LLM to strengthen the callee’s contract; in the latter, the system proceeds with further proof-level decomposition within the current method.
@@
3.3. Proof-level decomposition
@@ 4. Evaluation
LLMs selection. We primarily use Claude Sonnet 3.7 (Anthropic, 2025), as the Claude series models have demonstrated superior performance in generating Dafny code (Loughridge et al., 2024). Claude Sonnet 3.7 was the latest available version when this work began, ensuring it was not exposed to our experiments or prompts. Additionally, we utilize OpenAI o4-mini (OpenAI, 2025) for checking code verifiability, as it excels in complex reasoning at the time of submission. All the models run with the following configuration: initial temperature of $0.5$, the maximum number of tokens is $4028$, and the timeout is $20$ s. We selected these values to strike a balance between creativity, concise responses, and reasonable generation time.
Evaluation configuration. To prevent infinite verification loops, we set Dafny’s verification timeout to $20$ seconds. Within Prometheus, each node of the tree is allowed a maximum of $t=10$ generation attempts before rolling back. The regeneration counter $s$ is set to $2$, and $\Delta{temp}$ to $0.3$. This means that each node has two generation attempts with temperatures of $0.5$ and $0.2$. Finally, we set a hard timeout of $500$ seconds on the generation process for both the baseline and Prometheus, after which the attempt is aborted.
@@ 4.1. Verification Performance (Q1)
Prometheus achieves higher success rate on tasks involving nested loops and complex reasoning. We begin by demonstrating Prometheus’s ability to tackle non-trivial verification tasks using the TitanBench benchmark. As shown in Table 1, our system successfully completes $19$ out of $22$ tasks, achieving a success rate of 86%, compared to the baseline’s 68%. Among the $22$ tasks, $12$ involve programs with at least one nested loop. Importantly, the four additional successful verifications by Prometheus come from these nested loop tasks. While the baseline solves $6$ out of $10$ such tasks, Prometheus successfully solves all of them. Upon reviewing the solutions provided by Prometheus for these four tasks, we found that verifying them requires either $(i)$ multiple invariants for each loop, or $(ii)$ additional helper lemmas beyond a single method.
@@ 4.3. Robustness across formal specifications (Q3)
Prometheus effectively handles specification-code misalignment, especially in the presence of nested loops. Table 3 shows the results of the experiment, with tasks categorized by the presence/absence of nested loops. Among the 13 tasks, the baseline successfully verifies the 4 simplest ones. Notably, these tasks do not involve any arithmetic operations over the input array, which significantly simplifies verification. Without any arithmetic aggregation across the elements of the array, the inner loop does not need to maintain or reason about an evolving state, making the proof much more tractable. However, the baseline struggles significantly with tasks that include nested loops and aggregate state, aligning with our observations in Sec. 2.1 and Sec. 2.4 that AI-based systems tend to fail when multiple reasoning steps are required and to deal with misaligned, yet more readable, formal specifications. Conversely, Prometheus substantially outperforms the baseline by verifying $9$ among the $13$ tasks, including those with nested loops. Prometheus’ code decomposition mechanism enables it to modularize loops effectively, while its proof decomposition component introduces the necessary helper lemmas. Together, these features allow Prometheus to verify $4$ tasks that involve nested loops. In the three failed cases, Prometheus succeeds in proposing helper lemmas to bridge the misalignment, but fails in proving them. These lemmas are not trivial and require more advanced reasoning. We leave addressing this challenge for future work.
@@
4.4. Scaling to Complex Programs (Q4)
@@
The selected examples offer a balanced and meaningful range of complexity, with task difficulty determined by the number of verification lines and the number of helper lemmas required in the manually verified code. As shown in Table 4, the length of the verification code (LOVC) ranges from $21$ to $487$ lines, and the number of helper lemmas ranges from $1$ to $11$. Since the algorithms are more challenging to verify, we increase the generation timeout to $1500$ seconds. Because the complexity of these verification tasks surpasses the reasoning capacity of today’s LLMs, we supply a proof outline to both the baseline and Prometheus. [^4] The textual proofs are available in the artifact. These outlines ensure that the verification process begins from a sound and purposeful foundation, and they highlight how human insights can effectively complement LLM-based verification.
Prometheus can prove algorithms that require >>200 lines of verification code. Looking at the results in Table 4, the baseline is able to verify only the two simplest tasks, while Prometheus successfully verifies $7$ tasks within five runs. Interestingly, Prometheus often provides more lemmas than the corresponding human-written solutions, meaning that the system tends to decompose proofs into smaller, more manageable steps, while humans can deal with complex reasoning within a single, more comprehensive lemma. In the unsolved Task #8, Prometheus is capable of solving partial proof obligations but fails to generate all of them. We note that Prometheus starts to fail more often as the proof context grows in complexity. Nevertheless, the partially verified program can still be manually inspected by the user, who may complete the verification starting from the well-defined, partially correct proof.